OpenStudy (anonymous):

Write the slope-intercept equation of the line perpendicular to y = 5/2 x - 2, which passes through the point (0, 2). A. y = 5x + 2 B. y = -2/5x + 2 C. y = 2x + 5 D. y = 1/2x + 5

OpenStudy (anonymous):

do you know slope intercept formula? first of all

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

are you familiar with x axis y axis

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cools lets start from there

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do you know what a slope (non math)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

isnt the slope formula y=mx+b? or is it y1+y2=.....cant remeber

OpenStudy (anonymous):

ok slope intercept formula is y = m x + b y = y value x = x value m = slope

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

o ya i fogot b = y intercept (ill get to that later) what slope is rise (change or y) y = ------------------- ( x ) run (change of x)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

 not to scale rise 2 --- (x ) run 1 in fraction form 2 -- = 2 1 y = 2x

OpenStudy (anonymous):

good so far?

OpenStudy (anonymous):

yep followin yuh like ah trail:)

OpenStudy (anonymous):

now standard slope intercept form y = mx + b i left out b = y intercept i left out in the beginning i am going back to in now if i put y = 2x + 3 the I draw just starts from 3

OpenStudy (anonymous):

okay still getting it:)

OpenStudy (anonymous):

 not to scale